# version 0.9

### v5

@@
Yann LeCun
To my family: Dana, Hadar, Yoav, Oded and Deluca,
@@
The primary distinction between online nonstochastic control and other frameworks is the objective. In optimal control, robust control, and other control methodologies that assume stochastic noise, the goal is to perform comparably to an offline optimal strategy. In online nonstochastic control, both the cost functions as well as the perturbations from the assumed dynamical model are chosen by an adversary. Thus, the optimal policy is not defined a priori. Rather, the goal is to attain low regret against the best policy in hindsight from a benchmark class of policies.
@@ Acknowledgments
Ever since then, the gem of an idea turned out into a complete theory which is now mainstream in machine learning and control, with numerous applications and refinements. Notable events that contributed to this theory include the following. Max Simchowitz, who was a graduate student then, visited Princeton during 2019-2020, where he collaborated with the group, and the theory was extended to partial observed systems. Paula Gradu, who was then an undergraduate, and Edgar Minasyan, while a grad student, contributed to the study of time-varying and nonlinear dynamical systems. Anirudha Majumdar joined for several important projects, notably for iterative linear control. Udaya Ghai and Arushi Gupta studied extending the GPC to model-free RL.
@@
This book was further refined through teaching. Elad Hazan gratefully acknowledges the numerous contributions and insights of the students of the course computational control theory delivered at Princeton during COVID on Zoom during 2020-2021. The text was further enhanced by students of theoretical machine learning taught at Princeton University in the fall of 2024-2025.
The following colleagues and students were particularly helpful, making insightful comments and in finding various typos and suggesting improvements: Gon Buzaglo, Anand Brahmbhatt, Sofiia Druchyna, and Jonathan Pillow (and many more we wish to announce very soon).
@@
=$\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}$ definition
@@
$|A|$ determinant of matrix $A$
${\bf Tr}(A)$ trace of matrix $A$
$A\succeq 0$ $A$ is positive semi-definite, $\forall v\ ,\ v^{\top}Av\geq 0$
$A\succeq B$ $A-B$ is positive semi-definite
@@
$Q:S\times A\to{\mathbb{R}}$ $Q$ function
@@
Part I Background in Control and RL
@@ 1.1 What is This Book About?
The answer we give is not to apply gradient descent to any existing algorithm. Rather, we rethink the paradigm of control from the start. What can be achieved in a given dynamical system, whether known or unknown, fully observed or not, linear or non-linear, and without assumptions on the noise sequence?
This is a deep and difficult question that we fall short of answering. However, we do propose a different way of answering this question: we define the control problem from the point of view of an online player in a repeated game. This gives rise to a simple but powerful framework called nonstochastic control theory. The framework is a natural fit to apply techniques from online convex optimization, resulting in new, efficient, gradient-based control methods.
@@ 1.2 The Origins of Control
The industrial revolution was in many ways shaped by the steam engine that came to prominence in the 18th century. One of its most critical components was introduced by the Scottish inventor James Watt: the centrifugal governor, which is amongst the earliest examples of a control mechanism. A major source of efficiency losses of the earlier designs was the difficulty in regulating the engine. Too much or too little steam could easily cause engine failure. The centrifugal governor was created to solve this problem.
@@
The impact of the centrifugal governor on the steam engine, and in turn on the industrial revolution, illustrates the importance of the field of control in engineering and the sciences. Since the 18th century, our world has changed quite a bit, and the importance of control has only grown.
@@ 1.3 Formalization and Examples of a Control Problem
Given a dynamical system with starting state $(\mathbf{x}_{1},\mathbf{y}_{1})$, iteratively construct controls $u_{t}$ to minimize a sequence of loss functions $c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$, $c_{t}:{\mathbb{R}}^{d_{y}+d_{u}}\to{\mathbb{R}}$. This is given by the following mathematical program
@@
$s.t.$ $\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g_{t}(\mathbf{x}_{t}).$
@@
Instead of postulating a hypothesis about how our world should behave, we proceed to give an intuitive example of a control problem. We can then reason about the different questions and motivate the nonstochastc control setting.
@@ 1.3.1 Example: Control of a Medical Ventilator
In the figure below we see an illustration of a simple controller, that we will soon describe, called the PID controller, as it performs on a ventilator and lung simulator.
@@ 1.4.1 The Bang-Bang Controller
Bang-bang control is one of the simplest control strategies. Consider a scalar dynamical system, where it is desirable to keep the state $x_{t}$ within a prespecified dynamic range $[x_{\min},x_{\max}]$ (which can vary over time) and where the scalar control input $u_{t}$ must at all times be within the interval $[u_{\min},u_{\max}]$, which is assumed to contain zero for simplicity. A bang-bang controller chooses the control input at time $t$ as
@@
A bang-bang controller saturates its inputs to extreme values in an attempt to confine the state of the system to a desirable range. For example, in the context of mechanical ventilation, one could choose $[p_{t}^{*}-\Delta p,p_{t}^{*}+\Delta p]$ as a reasonable region around the prescribed pressure $p^{*}_{t}$ as a desirable range for $p_{t}$. In such a case, a bang-bang controller would allow oxygen to enter at the highest possible rate whenever the pressure drops below $p^{*}_{t}-\Delta p$.
@@ 1.4.2 The PID Controller
A more sophisticated control algorithm is the PID control method. Consider a control method that chooses the control to be a linear function of the observed states, i.e.
@@ 1.5 Classical Theory: Optimal and Robust Control
Optimal control postulates a probabilistic model for the perturbations, which are typically i.i.d. draws from a Gaussian distribution. However, other types of distribution can also be considered. As such, the control problem can be re-phrased in a more precise stochastic mathematical program,
@@
$s.t.$ $\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g(\mathbf{x}_{t}).$
@@
$s.t.$ $\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g(\mathbf{x}_{t}).$
The two formulations have been debated for decades, with pros and cons for each side. Optimal control focuses on the average case and often yields simpler formulations for optimization, resulting in more efficient methods. Robust control, on the other hand, allows for more difficult noise models and some amount of model misspecification. However, it is too pessimistic for many natural problems.
@@ 1.6 The Need for a New Theory
We would like an adaptive control theory which allows us to attain the best of both worlds: an efficient and fast flight when the weather permits, and a careful drone maneuver when this is required by the conditions. Can we design a control theory that will allow us to take into account the specific instance perturbations and misspecifications, and give us finite time provable guarantees? This is the subject of online non-stochastic control!
@@ 1.6.1 Online Nonstochastic Control Theory
This book is concerned with nonstochastic noise. When dealing with non-stochastic, i.e. arbitrary or even adversarial perturbations, the optimal policy, i.e. a decision making rule mapping observations to control inputs, is not clear a priori. Rather, an optimal policy for the observed perturbations can be determined only in hindsight. This is a significant difference from optimal and robust control theory, where the mapping from states to actions can be completely predetermined based upon the noise model.
Since optimality is defined only in hindsight, the goal is not predetermined. We thus shift to a different performance metric borrowed from the world of learning in repeated games, namely regret. Specifically, we consider regret with respect to a reference class of policies. In online control, the controller iteratively chooses a control $\mathbf{u}_{t}$. The controller then observes the next state of the system $\mathbf{x}_{t+1}$ and suffers a loss of $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ according to an adversarially chosen loss function. For simplicity, we consider the case of full state observation where $\mathbf{y}_{t}=\mathbf{x}_{t}$. Let $\Pi=\{\pi:\mathbf{x}\to\mathbf{u}\}$ be a class of policies. The regret of the controller with respect to $\Pi$ is defined as follows.
@@
where $\mathbf{u}_{t}={\mathcal{A}}(\mathbf{x}_{t})$ are the controls generated by ${\mathcal{A}}$, and $\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi}$ are the counterfactual state sequence and controls under the policy $\pi$, i.e.
@@
    First, the best policy in hindsight out of all possible policies may be very complicated to describe and to reason about. This argument was made by the mathematician Tyrell Rockafellar, see the bibliographic section for more details.
@@
The non-stochastic control problem can now be stated as finding an efficient algorithm that minimizes the worst-case regret vs. meaningful policies classes that we examine henceforth. It is important to mention that the algorithm does not have to belong to the comparator set of policies! Indeed, in the most powerful results, we will see that the algorithm will learn and operate over a policy class that is strictly larger than the comparator class.
@@ 1.6.2 A New Family of Algorithms
The policy regret minimization viewpoint in control leads to a new approach for algorithm design. Instead of a priori computation of the optimal controls, online convex optimization suggests modifying the controller according to the costs and dynamics, in the flavor of adaptive control, in order to achieve provable bounds on the policy regret.
@@
3:      Use control $\mbox{Use control }\mathbf{u}_{t}=\sum_{i=1}^{h}{{M}}_{i}^{t}\mathbf{w}_{t-i}$
@@
The version given here assumes that the time-invariant dynamics $f$ is known. The GPC computes a control that is a linear function of past perturbations. It modifies the linear parameters according to the loss function via an iterative gradient method. In Chapter 7, we prove that this algorithm attains sublinear regret bounds for a large class of time-varying and time-invariant dynamical systems. In later chapters, we explore extensions to nonlinear dynamical systems and unknown systems, which lie at the heart of nonstochastic control theory.
@@ 1.7 Bibliographic Remarks
There are many excellent textbooks on control theory, which we point to in later chapters. The book of [Ber07] offers extensive coverage of topics from the point of view of dynamic programming. The texts [Ste94, ZDG96a], on the other hand, delve deep into results on linear control, the most well-understood branch of control theory. Rockafellar [Roc87] proposed the use of convex cost functions to model the cost in a linear dynamical system to incorporate constraints on state and control and suggested computational difficulties with this approach. The text [SL91] gives a thorough introduction to adaptive control theory.
@@ 2.1.1 Medical Ventilator
where $\Delta$ is the time elapsed between successive actions of the discrete-time control algorithm. Being difficult to directly measure or estimate, the quantity $v_{t}$ is a hidden or latent variable from the controller’s perspective. Instead, the controller observes the air pressure $p_{t}$. We give a simplified physical model that relates the pressure and the volume. This model, given in the following, can be derived by first principles from stress-strain relations, treating the left and right lungs connected by the trachea as situated in a two-balloon experiment.
@@ 2.1.2 Double Integrator
The double integrator is a common example of a linear dynamical system. In its simplest form, it models a one-dimensional point object moving according to Newtonian mechanics on a one-dimensional trajectory (i.e. the real number line). The state $\mathbf{x}_{t}$ is 2-dimensional, with the first coordinate indicating the position of a point object and the second denoting its velocity. The control algorithm decides the force $u_{t}\in{\mathbb{R}}$ that is applied to the object at each time step. The dynamics equation can be written as
@@
Figure 2.2: The position and velocity visualized for a double integrator.
@@ 2.1.3 Pendulum Swing-up
where $\Delta$ is the time resolution of the controller. Note that this dynamical equation is non-linear in the state $\mathbf{x}_{t}=[\theta_{t},\dot{\theta_{t}}]^{\top}$ due to the presence of the term $\sin\theta$ in the evolution of $\dot{\theta_{t}}$.
Figure 2.3: A pivoted massless rod connected to a point object with mass $m$. The point object is to subject to gravitational acceleration $g$. Illustration by Krishnavedala, licensed under CC-BY-SA-3.0.

### v6

# version 1.0

@@
As someone who has worked extensively on learning theory and online learning, and later applied these ideas in domains such as autonomous driving and humanoid robotics, I find this book both timely and inspiring. It introduces a regret-minimization framework for control that draws on the elegance and power of online learning. Traditional control theory often models noise either as stochastic—sometimes unrealistically optimistic—or adversarial—often overly conservative. This book charts a new path by asking a deeper question: while we cannot predict noise, can we perform nearly as well as if we could? The answer, developed here, is a novel and exciting paradigm that bridges learning theory and control, and I believe it will have a lasting impact on both research and practice.
Shai Shalev-Shwartz
To my family: Dana, Hadar, Yoav, Oded, and Deluca,
@@
The primary distinction between online nonstochastic control and other frameworks is the objective. In optimal control, robust control, and other control methodologies that assume stochastic noise, the goal is to perform comparably to an offline optimal strategy. In online nonstochastic control, both the cost functions and the perturbations from the assumed dynamical model are chosen by an adversary. Thus, the optimal policy is not defined a priori. Rather, the goal is to attain low regret against the best policy in hindsight from a benchmark class of policies.
@@ Acknowledgments
Ever since then, the gem of an idea turned out into a complete theory that is now mainstream in machine learning and control, with numerous applications and refinements. Notable events that contributed to this theory include the following. Max Simchowitz, who was a graduate student then, visited Princeton during 2019-2020, where he collaborated with the group, and the theory was extended to partial observed systems. Paula Gradu, who was then an undergraduate, and Edgar Minasyan, while a graduate student, contributed to the study of time-varying and nonlinear dynamical systems. Anirudha Majumdar joined for several important projects, notably for iterative linear control. Udaya Ghai and Arushi Gupta studied extending the GPC to model-free RL.
@@
This book was further refined through teaching. Elad Hazan gratefully acknowledges the numerous contributions and insights of the students of the course computational control theory delivered at Princeton during COVID over Zoom during 2020-2021. The text was further enhanced by students of theoretical machine learning taught at Princeton University in the fall of 2024-2025.
The following colleagues and students were particularly helpful, making insightful comments, finding various typos and suggesting improvements: Gon Buzaglo, Anand Brahmbhatt, Sofiia Druchyna, Jonathan Pillow, and many more.
@@
$\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}$ definition
@@
$|A|$ determinant of matrix AA
${\bf Tr}(A)$ trace of matrix AA
$\succeq 0 AA$ is positive semi-definite, $\forall v\ ,\ v^{\top}Av\geq 0$
$A\succeq B$ ABA-B is positive semi-definite
@@
$S\times A\to{\mathbb{R}} QQ$ function
@@
Part I Background in Control and Reinforcement Learning
@@ 1.1 What is This Book About?
The answer we give is not to apply gradient descent to any existing algorithm. Rather, we rethink the paradigm of control from the start. What can be achieved in a given dynamical system, whether known or unknown, fully observed or not, linear or nonlinear, and without assumptions on the noise sequence?
This is a deep and difficult question that we fall short of answering. However, we do propose a different way of answering this question: we define the control problem from the point of view of an online player in a repeated game. This gives rise to a simple but powerful framework called nonstochastic control theory. The framework is a natural fit to applying techniques from online convex optimization, resulting in new, efficient, gradient-based control methods.
@@ 1.2 The Origins of Control
The industrial revolution was in many ways shaped by the steam engine that came to prominence in the eighteenth century. One of its most critical components was introduced by the Scottish inventor James Watt: the centrifugal governor, which is amongst the earliest examples of a control mechanism. A major source of efficiency losses of the earlier designs was the difficulty in regulating the engine. Too much or too little steam could easily cause engine failure. The centrifugal governor was created to solve this problem.
@@
The impact of the centrifugal governor on the steam engine, and in turn on the industrial revolution, illustrates the importance of the field of control in engineering and the sciences. Since the eighteenth century, our world has changed quite a bit, and the importance of control has only grown.
@@ 1.3 Formalization and Examples of a Control Problem
Given a dynamical system with starting state $(\mathbf{x}_{1},\mathbf{y}_{1})$, iteratively construct controls $u_{t}$ to minimize a sequence of loss functions $c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$, $c_{t}\colon{\mathbb{R}}^{d_{y}+d_{u}}\to{\mathbb{R}}$. This is given by the following mathematical program:
@@
$s.t.\ \$$\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g_{t}(\mathbf{x}_{t}).$
@@
Instead of postulating a hypothesis about how our world should behave, we proceed to give an intuitive example of a control problem. We can then reason about the different questions and motivate the nonstochastic control setting.
@@ 1.3.1 Example: Control of a Medical Ventilator
In figure 1.2 we see an illustration of a simple controller, which we will soon describe, called the PID controller, as it performs on a ventilator and lung simulator.
@@ 1.4.1 The Bang-Bang Controller
Bang-bang control is one of the simplest control strategies. Consider a scalar dynamical system, where it is desirable to keep the state $x_{t}$ within a prespecified dynamic range $x_{\min},x_{\max}]$ (which can vary over time) and where the scalar control input $u_{t}$ must at all times be within the interval $u_{\min},u_{\max}]$, which is assumed to contain zero for simplicity. A bang-bang controller chooses the control input at time $t$ as
@@
A bang-bang controller saturates its inputs to extreme values in an attempt to confine the state of the system to a desirable range. For example, in the context of mechanical ventilation, one could choose $p_{t}^{*}-\Delta p,p_{t}^{*}+\Delta p]$ as a reasonable region around the prescribed pressure $p^{*}_{t}$ as a desirable range for $p_{t}$. In such a case, a bang-bang controller would allow oxygen to enter at the highest possible rate whenever the pressure drops below $p^{*}_{t}-\Delta p$.
@@ 1.4.2 The PID Controller
A more sophisticated control algorithm is the PID control method. Consider a control method that chooses the control to be a linear function of the observed states, that is:
@@ 1.5 Classical Theory: Optimal and Robust Control
Optimal control postulates a probabilistic model for the perturbations, which are typically i.i.d. draws from a Gaussian distribution. However, other types of distribution can also be considered. As such, the control problem can be rephrased in a more precise stochastic mathematical program,
@@
$s.t.\ \$$\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g(\mathbf{x}_{t}).$
@@
$s.t.\ \$$\mathbf{x}_{t+1}=f_{t}(\mathbf{x}_{t},\mathbf{u}_{t},\mathbf{w}_{t}),\quad\mathbf{y}_{t}=g(\mathbf{x}_{t}).$
The two formulations have been debated for decades, with pros and cons for each side. Optimal control focuses on the average case and often yields simpler formulations for optimization, resulting in more efficient methods. Robust control, in contrast, allows for more difficult noise models and some amount of model misspecification. However, it is too pessimistic for many natural problems.
@@ 1.6 The Need for a New Theory
We would like an adaptive control theory that allows us to attain the best of both worlds: an efficient and fast flight when the weather permits, and a careful drone maneuvering when this is required by the conditions. Can we design a control theory that will allow us to take into account the specific instance perturbations and misspecifications, and give us finite-time provable guarantees? This is the subject of online nonstochastic control!
@@ 1.6.1 Online Nonstochastic Control Theory
This book is concerned with nonstochastic noise. When dealing with non-stochastic, that is, arbitrary or even adversarial perturbations, the optimal policy, that is, a decision making rule mapping observations to control inputs, is not clear a priori. Rather, an optimal policy for the observed perturbations can be determined only in hindsight. This is a significant difference from optimal and robust control theory, where the mapping from states to actions can be completely predetermined based upon the noise model.
Since optimality is defined only in hindsight, the goal is not predetermined. We thus shift to a different performance metric borrowed from the world of learning in repeated games, namely regret. Specifically, we consider regret with respect to a reference class of policies. In online control, the controller iteratively chooses a control $\mathbf{u}_{t}$. The controller then observes the next state of the system $\mathbf{x}_{t+1}$ and suffers a loss of $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ according to an adversarially chosen loss function. For simplicity, we consider the case of full state observation where $\mathbf{y}_{t}=\mathbf{x}_{t}$. Let $\Pi=\{\pi\colon\mathbf{x}\to\mathbf{u}\}$ be a class of policies. The regret of the controller with respect to $\Pi$ is defined as follows.
@@
where $\mathbf{u}_{t}={\mathcal{A}}(\mathbf{x}_{t})$ are the controls generated by $\mathcal{A}$}, and $\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi}$ are the counterfactual state sequence and controls under the policy $\pi$, that is,
@@
    First, the best policy in hindsight out of all possible policies may be very complicated to describe and to reason about. This argument was made by the mathematician Tyrell Rockafellar; see the bibliographic section for more details.
@@
The non-stochastic control problem can now be stated as finding an efficient algorithm that minimizes the worst-case regret vs. meaningful policy classes that we examine henceforth. It is important to mention that the algorithm does not have to belong to the comparator set of policies! Indeed, in the most powerful results, we will see that the algorithm will learn and operate over a policy class that is strictly larger than the comparator class.
@@ 1.6.2 A New Family of Algorithms
The policy regret minimization viewpoint in control leads to a new approach to algorithm design. Instead of a priori computation of the optimal controls, online convex optimization suggests modifying the controller according to the costs and dynamics, in the flavor of adaptive control, in order to achieve provable bounds on the policy regret.
@@
3:   $\mbox{Use control }\mathbf{u}_{t}=\sum_{i=1}^{h}{{M}}_{i}^{t}\mathbf{w}_{t-i}$
@@
The version given here assumes that the time-invariant dynamics $f$ are known. The GPC computes a control that is a linear function of past perturbations. It modifies the linear parameters according to the loss function via an iterative gradient method. In Chapter 7, we prove that this algorithm attains sublinear regret bounds for a large class of time-varying and time-invariant dynamical systems. In later chapters, we explore extensions to nonlinear dynamical systems and unknown systems, which lie at the heart of nonstochastic control theory.
@@ 1.7 Bibliographic Remarks
There are many excellent textbooks on control theory, which we point to in later chapters. The book of [Ber07] offers extensive coverage of topics from the point of view of dynamic programming. The texts [Ste94, ZDG96a], in contrast, delve deeply into results on linear control, the most well-understood branch of control theory. Rockafellar [Roc87] proposed the use of convex cost functions to model the cost in a linear dynamical system to incorporate constraints on state and control and suggested computational difficulties with this approach. The text [SL91] gives a thorough introduction to adaptive control theory.
@@ 2.1.1 Medical Ventilator
where $\Delta$ is the time elapsed between successive actions of the discrete-time control algorithm. Being difficult to directly measure or estimate, the quantity $v_{t}$ is a hidden or latent variable from the controller’s perspective. Instead, the controller observes the air pressure $p_{t}$. We give a simplified physical model that relates the pressure and the volume. This model, given in the following, can be derived by first principles from stress-strain relations, treating the left and right lungs connected by the trachea as situated in a two-balloon experiment:
@@ 2.1.2 Double Integrator
The double integrator is a common example of a linear dynamical system. In its simplest form, it models a one-dimensional point object moving according to Newtonian mechanics on a one-dimensional trajectory (i.e. the real number line). The state $\mathbf{x}_{t}$ is two dimensional, with the first coordinate indicating the position of a point object and the second denoting its velocity. The control algorithm decides the force $u_{t}\in{\mathbb{R}}$ that is applied to the object at each time step. The dynamics equation can be written as
@@
[…]
Figure 2.2: Double integrator illustration showing state coordinate $x_{1}(t)$ (position), state coordinate $x_{2}(t)$ (velocity), mass $m$, and control input $u(t)$.
@@ 2.1.3 Pendulum Swing-Up
where $\Delta$ is the time resolution of the controller. Note that this dynamical equation is nonlinear in the state $\mathbf{x}_{t}=[\theta_{t},\dot{\theta_{t}}]^{\top}$ due to the presence of the term $\sin\theta$ in the evolution of $\dot{\theta_{t}}$.
PivotmmAngle $\theta_{t}$$m\mathbf{g}$$\mathbf{u}_{t}$$l$
Figure 2.3: Pendulum swing-up illustration with marked angle $\theta_{t}$, gravitational force mgmg, control input torque $u_{t}$, mass $m$, and rod length $l$.